Redford Company had January 1 inventory of $300,000 when it adopted dollar-value LIFO. During the year, purchases were $1,800,000 and sales were $3,000,000. December 31 inventory at year-end prices was $430,080, and the price index was 115. How much is Redford's ending inventory at December 31? NOTE: Show your work and label your numbers. Round all calculations to whole dollars.
Answer #1

(A) January 1 Inventory = $300000

(B) December 31 Inventory = $430080

(C) Price Index = 115

(D) Dec 31 Inventory Price according to Base Year = B/C * 100

$430080/115 * 100 = $373983

(E) = D - A

= $373983 - $300000 = $73983

(F) = E X C/100

= $73983 * 115/100 = $85080

$300000 + $85080 = $385080